Circular No. 6203

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) C. Lidman and P. Bouchet, European Southern Observatory (ESO),

report observations of C/1995 O1 in the near infrared with the 2.2-m telescope at La Silla (+ ESO IRAC2 camera with Nicmos3 array of 256×256 pixels; pixel size 0′′.52): “The observations were centered at Aug. 5.14 UT. The following preliminary magnitudes (± 0.1 mag) were obtained through a 10′′ aperture centered on the comet’s nucleus: J = 11.2, H =11.2, K = 11.1. The images show a clear extension of the order of 1′ toward the northwest, the coma size being ∼ 2′′.5 (seeing 0′′.8). Notable is that the comet is colorless in the near-infrared.”

SUPERNOVA 1995V IN NGC 1087 D. Balam, Climenhaga Observatory, reports the following precise po-

sition, obtained with the 0.5-m reflector (+ CCD): α = 2^{h}46^{m}26^{s}.77, δ = −0^{o}29′55′′.6 (equinox 2000.0), which is 23′′ east and 0′′.2 south of the host galaxy’s core. Preliminary photometry in Johnson V and Cousins I bands: Aug. 4.47 UT, V = 15.68 ± 0.09, V − I = +1.53 ± 0.09 (comparison stars HD 18175, HD 18369). A nearby star of similar brightness (V = 14.94 ± 0.04, V − I = +0.30 ± 0.04) has position end figures 26^{s}.75, 25′37′′.4.

X-RAY NOVA 1994 IN SCORPIUS J. Orosz, B. Schaefer, and S. Barnes, Yale University, report: “We have

observed the superluminous γ-ray source GRO J1655–40 with the Cerro Tololo 0.9-m telescope (+ CFCCD imager) from June 26 to July 3 and from Aug. 4 to 7 UT. The source is currently almost 1 mag brighter in the V band than it was during the period Apr. 5–24. The average magnitude was V = 15.9 during each of the two recent runs, compared to the average of V = 16.7 during April. This increase in brightness may be related to the recent increase in the hard x-ray activity (IAUC 6196 and 6201), but it should be noted that our observations in Mar./Apr. (IAUC 6173) showed no response to the hard x-ray outburst observed at that time (IAUC 6152). The mean V − I color was +1.93 during each of the two recent runs, a decrease of ∼ 0.05 mag since Apr. The June/July V light curve, folded on the 2.6-day period (IAUC 6173), has a similar shape as the folded Apr. V light curve, but with half the amplitude. Also, the modulation in V − I in June/July is only 0.04 mag, compared with 0.15 mag in Apr. Additional observations are encouraged.”

1995 August 8 Daniel W. E. Green